OpenStudy (anonymous):

For people smarter than me, plz help! A graduate student is comparing final-exam test scores of male and female students in an introductory physics class. She starts with no expectations as to which sex will score more highly. After seeing that men did better than women on the first quiz, she tests a one-sided alternative about the mean final-exam scores, She finds z=1.9 with one-sided P- value is P=0.0272. She should have used the two-sided alternative hypothesis. one-sided alternative hypothesis. What is the correct P-value (0.0001) for z=1.9?

OpenStudy (anonymous):

p-value applicable to one tailed test = P(Z > z) = P(Z > 1.9) = 0.0287 p-value applicable to two tailed test = 2*P(Z > z) = 2*P(Z > 1.9) = 2*0.0287 = 0.0574 One sided alternative hypothesis is to be used Correct p-value for z = 1.9 is 0.0287

OpenStudy (anonymous):

so yes it would be the correct answer

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.